# A Practical Framework for Estimating the Repetition Likelihood of Fast Radio Bursts from Spectral Morphology

### v1

@@ 4 Conclusion
    We construct a four-quadrant probability map in the $\gamma$-$r$ space to rapidly assess the likelihood of repetition for different spectral parameter combinations. This map exhibits a pronounced statistical gradient, with repetition likelihoods of $\sim$65% in the high-repetition region and only $\sim$5% in the low-repetition region, demonstrating that the spectral morphology can serve as an effective predictor of repeatability tendencies. Notably, all three bursts from FRB 20201124A fall within the high-probability Region I, validating the robustness of this framework. Thus, it provides a simple, quantifiable, and easily generalizable tool for predicting repetition probabilities, serving as a practical reference for future telescope observations and studies of FRB repeatability. In addition, we identify 40 apparent non-repeaters with probability scores above 0.6 in Region I and recommend them as priority targets for future monitoring (details in the Appendix).
@@ Acknowledgments
We are very grateful to Alice P. Curtin, Ziggy Pleunis, and Mary Jiang for helpful discussions. We acknowledge the support of the National SKA Program of China (Grant Nos. 2022SKA0110200, 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 12533001, 12575049, and 12473001), the China Manned Space Program (Grant No. CMS-CSST-2025-A02), and the National 111 Project (Grant No. B16009). F.-W.Z. acknowledges the supports from the National Natural Science Foundation of China (No. 12463008) and the Guangxi Natural Science Foundation (No. 2022GXNSFDA035083).
@@
Note. — The ‘Sub_num’ column represents the sub-burst number assigned to each FRB in the CHIME/FRB sample. The ‘Quadrant’ column denotes the $\gamma$-$r$ quadrant to which each FRB belongs (see Section 3.4 for details).

### v2

@@ IV Conclusion
    We construct a four-quadrant probability map in the $\gamma$-$r$ space to rapidly assess the likelihood of repetition for different spectral parameter combinations. This map exhibits a pronounced statistical gradient, with repetition likelihoods of $\sim$67% in the high-repetition region and only $\sim$7% in the low-repetition region, demonstrating that the spectral morphology can serve as an effective predictor of repeatability tendencies. Notably, all three bursts from FRB 20201124A fall within the high-probability Region I, validating the robustness of this framework. Thus, it provides a simple, quantifiable, and easily generalizable tool for predicting repetition probabilities, serving as a practical reference for future telescope observations and studies of FRB repeatability. In addition, we identify 40 apparent non-repeaters with probability scores above 0.6 in Region I and recommend them as priority targets for future monitoring (details in the Appendix).
@@ Acknowledgments
We are very grateful to Alice P. Curtin, Ziggy Pleunis, and Mary Jiang for helpful discussions. We acknowledge the support of the National Natural Science Foundation of China (Grants Nos. 12473001, 12533001, and 12575049), the National SKA Program of China (Grant Nos. 2022SKA0110200, 2022SKA0110203), the China Manned Space Program (Grant No. CMS-CSST-2025-A02), and the 111 Project (Grant No. B16009). F.-W.Z. acknowledges the supports from the National Natural Science Foundation of China (No. 12463008) and the Guangxi Natural Science Foundation (No. 2022GXNSFDA035083). Y.-K.Z. is supported by the Postdoctoral Fellowship Program and China Postdoctoral Science Foundation (Grant Number BX20250158).
@@
Note. — The ‘Sub_num’ column represents the sub-burst number assigned to each FRB in the CHIME/FRB sample. The ‘Quadrant’ column denotes the $\gamma$-$r$ quadrant to which each FRB belongs (see Section III.4 for details).